OpenStudy (anonymous):

if sin(x-pi/2)=.45 find cos(-x) would it just be -.45?

jimthompson5910 (jim_thompson5910):

cos(-x) = cos(x) cos(-x) = sin(pi/2 - x) cos(-x) = sin(-1(-pi/2 + x)) cos(-x) = sin(-1(x - pi/2)) cos(-x) = -sin(x - pi/2) cos(-x) = -(0.45) cos(-x) = -0.45 So you are correct. Nice work.

OpenStudy (anonymous):

@jim_thompson5910 really yay

jimthompson5910 (jim_thompson5910):

yes, you nailed it kkayla

OpenStudy (anonymous):

@jim_thompson5910 :)